# Explainable Malware Analysis: Concepts, Approaches and Challenges

### v1

@@ Abstract
Machine learning (ML) has seen exponential growth in recent years, finding applications in various domains such as finance, medicine, and cybersecurity. Malware remains a significant threat to modern computing, frequently used by attackers to compromise systems. While numerous machine learning-based approaches for malware detection achieve high performance, they often lack transparency and fail to explain their predictions. This is a critical drawback in malware analysis, where understanding the rationale behind detections is essential for security analysts to verify and disseminate information. Explainable AI (XAI) addresses this issue by maintaining high accuracy while producing models that provide clear, understandable explanations for their decisions. In this survey, we comprehensively review the current state-of-the-art ML-based malware detection techniques and popular XAI approaches. Additionally, we discuss research implementations and the challenges of explainable malware analysis. This theoretical survey serves as an entry point for researchers interested in XAI applications in malware detection. By analyzing recent advancements in explainable malware analysis, we offer a broad overview of the progress in this field, positioning our work as the first to extensively cover XAI methods for malware classification and detection.
@@
Explainable malware analysis, Interpretable malware analysis, Malware classification, and Malware detection.
@@ I Introduction
In today’s digital era, malware poses a formidable threat, causing significant damage to computer systems and resulting in billions in financial losses. This pervasive issue has economic impacts and restricts users’ access to essential services worldwide. Especially with the rise of sophisticated malware such as zero-day malware, traditional detection methods are becoming increasingly inadequate. Hence, more advanced and automated malware detection solutions should be considered.
Recently, Machine Learning (ML) and Deep Learning (DL) techniques has shown the ability to detect known and zero-day malware. However, to be able to detect malware with high accuracy, traditional ML models tend to become very sophisticated and large, making prediction interpretability a difficult task. Similarly, DL models are inherently sophisticated (often labeled as black-box models) as their internal decision-making processes are not easily interpretable. This opaqueness can be a significant issue, especially when the rationale behind predictions remains unclear [96].
This lack of interpretability is particularly concerning for cybersecurity professionals and end-users who rely on these ML-based malware detectors for protection since they require clear explanations behind the predictions. Without explainability, it is challenging to trust the performance of these systems to identify and correct errors in detection, analyze emerging or zero-day threats, and ensure fairness and regulatory compliance. To address this, explainable AI (XAI) has seen a growing interest. It provides a method for making decisions or predictions that are accurate and transparent and offers users a clear and comprehensible explanation of the reasoning behind the conclusions [100].
Several survey papers [32], [124], [51], [121], [21] have focused on XAI and conducted comprehensive reviews on the research areas, methods, and opportunities of explainable ML models, offering both mathematical and visual explanations. In research done by Danilevsky et al. [31], they focus on a survey of XAI in NLP and its evaluation in that area. Another study [84] presents a framework for the evaluation of this approach along with a systematic survey on XAI. Another research [66] centers on knowledge-driven and data-driven methods in XAI, whereas [29] discusses the historical development of XAI and its application in expert systems. Speith [110] offers an overview of the general taxonomy of XAI approaches and the challenges in this field and extends the discussion to include challenges faced by researchers.
Saeed and Omlin [97] conduct a systematic literature review related to challenges and research directions in the XAI field. It provides information regarding general challenges on XAI, and then it focuses on challenges and potential research directions based on the phases of the ML lifecycle—design, development, and deployment.
The systematic review by Saranya and Subhashini [100] provides a thorough overview of XAI developments in various fields such as agriculture, social media, computer vision, and healthcare. This review also discusses the challenges in these areas. Milani et al. [82] provide a survey on explainable reinforcement learning, which is used to show the decision-making process for reinforcement learning agents. They overview the literature on explainable reinforcement learning and show the future directions in this scope. Nasser and Nasser [87] provide a survey of the application of hardware capabilities and ML to enhance malware detection. They explore various hardware architectures combined with ML models for dynamic and efficient malware identification. The paper also discusses the potential future developments in hardware-assisted XAI for malware detection.
Charmet et al. [22] study the intersection of XAI and cybersecurity with a focus on two main objectives: exploring the application of XAI to enhance cybersecurity defenses and also the security of XAI methods against potential adversarial attacks. This survey contributes to integrating XAI into cybersecurity practices to have robust, understandable, and trustworthy AI deployments. However, as the research covers the whole concept of cybersecurity, it does not consider various studies on explainable ML in the context of malware analysis.
These surveys, published between 2019 and 2024, highlight the increasing interest in XAI across various domains, including medical, NLP, and cybersecurity. Despite this extensive body of work, there remains a noticeable gap. There is no survey that exclusively addresses the malware domain within the realm of XAI. Moreover, while the terms ’interpretable’ and ’explainable’ AI are often used interchangeably, there is a distinct lack of differentiation in the existing literature. An exception is the work by Lin and Chang, [68], which specifically examines the taxonomy of interpretable malware detectors for ML-based models, focusing on interpretable malware detection. In contrast, our survey broadens the scope to contain both interpretable and explainable methods in malware classification and detection. We introduce a novel taxonomy for explainability approaches in this sector and tabulate the latest advancements in explainable malware analysis based on this taxonomy. The limited research focusing on malware in the context of XAI made us conduct this survey, which thoroughly covers explainable malware analysis, its approaches, and the open research challenges in the field. Hence, this paper contributes significantly to the field of XAI with a particular focus on malware detection. The contributions of this paper are as follows:
@@
  • Our work presents an extensive survey covering various XAI models and techniques used across multiple disciplines. This contribution offers a broad view of XAI, showcasing its applications and relevance in different areas.
@@
  • We provide an in-depth overview of ML-based approaches in malware detection to understand the intersection of ML and malware detection, which fills a notable gap in the current literature.
@@
The structure of the remainder of this paper is outlined as follows: Section II offers an in-depth exploration of file classification and online malware detection methods. Section III discusses ML-based models and the explainable techniques. Section IV is dedicated to the studies on approaches and techniques in explainable malware classification and detection. This is followed by Section V, which addresses the open challenges and future prospects in this area. Finally, The paper concludes with Section conclusion, which provides a summary of our work.
@@ II Malware Detection Approaches
Malware detection techniques are used to address the threat posed by malware. They are generally categorized into two distinct approaches: File classification and Online-based approaches. This categorization is clearly illustrated in Figure 1. The field has seen considerable research efforts, with numerous studies and developments aimed at enhancing the efficacy and reliability of these malware detection methodologies.
Figure 1: Malware Detection Approaches
@@ II-A File Classification Approach
File classification focuses on the analysis of a file’s code to determine whether it is malware. The process begins with the identification of a potentially suspicious file. To thoroughly assess its nature, file classification employs diverse methods, which fall into three main categories: Static analysis, Dynamic analysis, and Hybrid analysis. Static analysis involves examining the file’s code without executing it and looking for malicious patterns. In contrast, in dynamic analysis, the file is executed in a secured environment to observe and analyze its behavior. Hybrid analysis combines these two approaches, leveraging the strengths of both static and dynamic examinations. Once a file is concluded to be non-malicious, it is generally exempt from ongoing scrutiny. These varying techniques in file classification are designed to address different aspects of malware detection.
@@ II-A1 Static Analysis
Static analysis involves the careful examination of an executable’s signature without the need to execute the code, aiming to classify the file as malware if the signature appears malicious or as benign if otherwise [30]. This method has the reverse engineering of malware code and involves the detailed processing of extracted features to discern and interpret any malicious activities through a signature-based approach. In this context, a signature refers to a unique identifier for a binary file, determined by calculating its cryptographic hash. Extensive research, notably by Hou et al.[48] and Kim and Lee [53], has been dedicated to enhancing static malware detection, with a particular focus on the extraction of Application programming interfaces (API) calls from Portable Executable (PE) files using techniques like stacked autoencoders. This process involves extracting vital features such as API calls, Opcode sequences, and N-Grams from potentially suspicious files as illustrated in Figure 2, which are then employed to train ML algorithms for more accurate and efficient malware detection. For instance, the work of Shankarapani et al. [103] has been using API and Opcode sequences to effectively identify segments of code that closely resemble known malware patterns. However, it is important to recognize that static analysis, while valuable, is not without its limitations. One significant challenge is its inability to detect malware that is actively running within a system or to identify completely new malware variants that have not been cataloged.
Figure 2: Static and Dynamic Malware Detection Approaches
@@ II-A2 Dynamic Analysis
It entails executing malware within a secure virtual environment, such as a cuckoo sandbox, to study its behavior meticulously [30] [52]. This method is particularly effective in addressing zero-day malware threats. The dynamic analysis process, as depicted in Figure 2, starts with executing a suspicious PE file in a sandbox environment, ensuring isolation from external systems. This controlled execution allows for the collection of essential data, including memory features, system calls, and function calls. Subsequently, these collected data are preprocessed and used to train various ML-based algorithms, which can enhance the malware detection model. Unlike static analysis, dynamic malware analysis requires the execution of code in a time-restricted, closed environment, which can be resource-intensive. Research endeavors, as demonstrated by Firdausi et al. [38] and Luckett et al. [72], have utilized system calls as key features for training traditional ML models like k-Nearest Neighbor (K-NN), Decision Tree, Support Vector Machine (SVM), and Naive Bayes. Furthermore, studies by Pirscoveanu et al. [91], and Tobiyama et al. [112] have focused on the extraction of features from API calls, evaluating the effectiveness of ML algorithms, including Random Forest, k-NN, and Convolutional Neural Networks (CNN) in dynamic malware analysis. These approaches highlight the dynamic method’s capacity for dealing with complex malware detection challenges, although it requires significant time and resource allocation.
@@ II-A3 Hybrid Analysis
A methodology integrating static and dynamic techniques, hybrid analysis, is another malware detection technique [114]. This concept has been explored in various studies. For example, Santoso et al. [99] utilized a combination of Artificial Neural Networks (ANN) and CNN for malware detection, yielding accuracies of 99.4% and 97%, respectively. Focusing on Android malware, Zhu et al.[130] proposed an innovative framework using the Merged Sparse Autoencoder (MSAE), which is an unsupervised learning algorithm demonstrating its effectiveness. Adding to this, Tong and Yan [113] developed a method that combines static and dynamic analysis for mobile malware detection. This method involves comparing system call patterns of benign and malicious applications with the dynamic analysis applied to unknown applications. Subsequent offline comparison of these pattern sets further validates the unknown application’s nature. Their results show the advantages of the hybrid approach over methods relying solely on static or dynamic analysis. Altaher and Barukab [7] also proposed a hybrid methodology for Android malware detection that leverages API calls and application permissions, further substantiating the potential of hybrid techniques in this field.
@@ II-B Online Malware Detection
Online malware detection stands out as a distinct approach in the cybersecurity domain. Unlike static, dynamic, or hybrid methods that analyze specific malware samples, online detection monitors the entire system in real-time, which enables the capture of malware at any moment, regardless of its activity level. This technique focuses on the behavior of the entire machine rather than individual malware behaviors.
Key contributions in this area include the work of Watson et al. [120], who developed a system using performance metrics to build SVM, achieving a 90% accuracy rate. Azmandian et al. [15] proposed intrusion-based detection techniques, while Abdelsalam et al. [2] introduced a sequential k-means clustering algorithm for anomaly detection, specifically designed for a standard 3-tier architecture on an OpenStack testbed. Their approach, notably effective for high-profile malware, leverages virtual machine systems and resource utilization features but shows limitations in detecting low-resource-utilization malware. Further research was done by McDole et al. [78], who examined various CNN models to determine their suitability for malware detection in cloud Infrastructure as a Service (IaaS). Their subsequent study [79] compared the process-level performance metrics of different deep learning models in the context of online malware detection in cloud IaaS environments. Similarly, Kimmel et al. [55, 54] presented a comprehensive analysis of the effectiveness of several ML models for online malware detection, focusing on system features describing processes in a virtual machine. They emphasized the use of CNNs, which are known for their simplicity and effective representation in 2D format. Abdelsalam et al. [1] extended this concept by employing a 3-dimensional CNN to enhance classifier accuracy and specifically target low-profile malware, achieving an accuracy rate of 90%.
@@ III Explainability in Machine Learning
In recent years, the fields of ML and AI have gained significant attention due to their potential in addressing complex real-world problems. These technologies, particularly ML models, have demonstrated a remarkable ability to make accurate predictions based on training data. However, an important challenge arises from the inherent complexity of these models, often making them difficult to interpret. This complexity frequently results in what is termed black-box models, where the internal workings and decision-making processes remain opaque and not easily understandable [46].
DL-based models enable machines to develop complex hierarchical data patterns, which play a key role in tasks like classification or detection. These models, by layering and integrating various levels of data representation, can enhance the predictive power of systems. However, this increased complexity often obscures the internal decision-making process, which can lead to questions about their decision logic [69, 11]. In contrast, white-box models offer a more transparent approach. They are designed to be easily interpretable, which allows users to understand how input data is transformed into predictions or decisions. This transparency is particularly valuable in fields where understanding the reasoning behind a decision is as important as the decision itself [47, 35].
@@
This is where the concept of XAI comes into the picture. As depicted in Figure 3, today’s AI systems typically involve training the data, undergoing the machine learning process, and providing prediction for end-users. In contrast, XAI goes a step further. It can deliver high-accuracy predictions and provides clear, justifiable explanations for these outcomes. This transparency in AI decision-making processes enhances user trust. With higher interpretability, the reasons behind AI predictions become more comprehensible to humans, which boosts the trustworthiness and reliability of the model’s predictions. In the comprehensive study, Blanco-Justicia and Domingo-Ferrer. [18] discussed the characteristics that define XAI for enhancing transparency and efficacy in AI systems as follows:
Figure 3: Explainable Machine Learning Concept
@@
Hence, in the context of XAI, it is important to understand the general classification of ML-based models, which is illustrated in Figure 4. This figure presents a comprehensive taxonomy of ML models and XAI techniques and provides a clear framework for understanding this field.
@@
ML models fall into two primary categories: Transparent and Opaque. Transparent models are inherently explainable. These models are straightforward enough that they do not require additional post-hoc explainability techniques, i.e., techniques provide explanations only after the training process has finished. However, when the processes within these models become more complex, there arises a need for post-hoc explainability to make their functioning clear.
On the other hand, Opaque models, often referred to as black-box models, are characterized by their high accuracy yet present significant challenges in interpretation. Due to their complexity, they require the use of post-hoc explainability methods. The goal of post-hoc explainability is to make the outcomes of ML models more transparent, understandable, and trustworthy to humans. Post-hoc explainability can be further divided into two types: model-agnostic and model-specific methods. Model-agnostic methods have a variety of explainability techniques and are versatile enough to be applied to any black-box model. These methods are compatible with a wide range of ML models and offer flexibility in interpretation. In contrast, model-specific methods are applicable only to certain types of models and limit their utility to specific cases.

### v2

@@ Abstract
Machine learning (ML) has rapidly advanced in recent years, revolutionizing fields such as finance, medicine, and cybersecurity. In malware detection, ML-based approaches have demonstrated high accuracy; however, their lack of transparency poses a significant challenge. Traditional black-box models often fail to provide interpretable justifications for their predictions, limiting their adoption in security-critical environments where understanding the reasoning behind a detection is essential for threat mitigation and response. Explainable AI (XAI) addresses this gap by enhancing model interpretability while maintaining strong detection capabilities. This survey presents a comprehensive review of state-of-the-art ML techniques for malware analysis, with a specific focus on explainability methods. We examine existing XAI frameworks, their application in malware classification and detection, and the challenges associated with making malware detection models more interpretable. Additionally, we explore recent advancements and highlight open research challenges in the field of explainable malware analysis. By providing a structured overview of XAI-driven malware detection approaches, this survey serves as a valuable resource for researchers and practitioners seeking to bridge the gap between ML performance and explainability in cybersecurity.
@@
Explainable Malware Analysis, Interpretable Malware Analysis, Explainable AI, AI for Security, Malware Detection, and Malware Classification.
[…]
=-15pt
@@ I Introduction
In today’s digital landscape, malware remains a formidable threat, causing billions in financial losses and disrupting critical services worldwide. The increasing sophistication of attacks—particularly zero-day malware—has rendered traditional detection and analysis methods increasingly ineffective. As a result, there is a need for advanced, automated malware detection solutions that can adapt to evolving threats.
Machine Learning (ML) and Deep Learning (DL) have emerged as powerful tools for malware detection, demonstrating the ability to identify both known and zero-day threats. However, to achieve high accuracy, these models often grow complex and opaque, making it difficult to understand how predictions are made. DL-based models, in particular, are frequently described as black boxes, as their decision-making processes remain largely inscrutable to users and security professionals alike [1]. This lack of interpretability poses a significant challenge in cybersecurity, where understanding why a detection occurred is just as important as the detection itself for ensuring reliability, fairness, and error analysis.
To address this issue, Explainable AI (XAI) has gained increasing attention. XAI aims to bridge the gap between accuracy and explainability by providing transparent, comprehensible explanations for model predictions. By making malware detection systems more explainable, XAI enhances trust, facilitates threat analysis, and enables security professionals to make informed decisions based on model outputs [2].
Several surveys [3, 4, 5, 6, 7] have reviewed XAI, covering research areas, methods, and opportunities with mathematical and visual explanations. Danilevsky et al. [8] focus on XAI in NLP, while Mohseni et al. [9] propose an evaluation framework. Li et al. [10] explore knowledge-driven and data-driven methods, and Confalonieri et al. [11] discuss XAI’s evolution in expert systems. Speith [12] provides a taxonomy of XAI approaches and challenges.
Saeed and Omlin [13] review XAI challenges and research directions, while Milani et al. [14] survey explainable reinforcement learning. Nasser and Nasser [15] examine hardware-assisted ML for malware detection. Charmet et al. [16] explore XAI’s role in cybersecurity but do not specifically address explainable ML in malware analysis.
Despite growing interest in XAI (2019–2024), no survey exclusively focuses on malware analysis. Existing literature lacks a clear distinction between ‘interpretable’ and ‘explainable’ AI, except for Lin and Chang [17], who categorize interpretable malware detectors. Our survey fills this gap by covering both interpretable and explainable methods in malware classification and detection, introducing a taxonomy for explainability approaches, and summarizing recent advancements. This work aims to provide a comprehensive view of explainable malware analysis, its methodologies, and open research challenges.
Hence, this paper contributes significantly to the field of XAI with a particular focus on malware analysis. The key contributions are as follows:
@@
  • Our work presents an extensive survey covering various XAI models and techniques used across multiple disciplines. This contribution offers a broad view of XAI and showcases its applications and relevance in different areas.
@@
  • We provide an in-depth overview of ML-based approaches in malware detection to understand the intersection of ML and malware detection, which fills a gap in the current literature.
@@
The structure of the remainder of this paper is outlined as follows: Section II offers an in-depth exploration of file classification and online malware detection methods. Section III discusses ML-based models and the explainable techniques. Section IV is dedicated to the studies on approaches and techniques in explainable malware classification and detection. This is followed by Section V, which addresses the open challenges and future prospects in this area. Finally, The paper concludes with a conclusion, which provides a summary of our work.
@@ II Malware Detection Approaches
Malware detection techniques are used to detect the threat posed by malware. They are generally categorized into two distinct approaches: File Classification and Online-Based Approaches. The field has seen considerable research efforts, with numerous studies and developments aimed at enhancing the efficacy and reliability of these malware detection methodologies.
Figure 1: Machine Learning-Based File Classification Techniques
@@ II-A File Classification Approach
File classification focuses on the analysis of a file’s code to determine whether it is malware. The process begins with the identification of a potentially suspicious file. To thoroughly assess its nature, file classification employs different methods, which fall into three main categories: Static analysis, Dynamic analysis, and Hybrid analysis.
Static analysis involves examining the file’s code without executing it and looking for malicious patterns. In contrast, in dynamic analysis, the file is executed in a secured environment to observe and analyze its behavior. Hybrid analysis combines these two approaches, leveraging the strengths of both static and dynamic examinations. Once a file is concluded to be non-malicious, it is generally exempt from ongoing scrutiny. These varying techniques in file classification are designed to address different aspects of malware detection. This categorization, along with the explanation of dynamic and static analysis processes, are also illustrated in Fig 1.
@@ II-A1 Static Analysis
Static analysis involves the careful examination of an executable’s signature without the need to execute the code, aiming to classify the file as malware if the signature appears malicious or as benign if otherwise [18]. This method has the reverse engineering of malware code and involves the detailed processing of extracted features to discern and interpret any malicious activities through a signature-based approach. In this context, a signature refers to a unique identifier for a binary file, determined by calculating its cryptographic hash.
Multiple research, e.g. studies by Hou et al.[19] and Kim and Lee [20], have been dedicated to enhancing static malware detection, with a particular focus on the extraction of Application programming interfaces (API) calls from Portable Executable (PE) files using techniques like stacked autoencoders. This process involves extracting vital features such as API calls, Opcode sequences, and N-Grams from potentially suspicious files, as illustrated in Fig 1, which are then employed to train ML algorithms for more accurate and efficient malware detection.
For instance, the work of Shankarapani et al. [21] has been using API and Opcode sequences to effectively identify segments of code that closely resemble known malware patterns. However, it is important to recognize that static analysis, while valuable, is not without its limitations. One significant challenge is its inability to detect malware that is actively running within a system or to identify completely new malware variants that have not been cataloged.
@@ II-A2 Dynamic Analysis
It involves executing malware within a secure virtual environment, such as a cuckoo sandbox, to study its behavior meticulously [18] [22]. This method is particularly effective in addressing zero-day malware threats. The dynamic analysis process, as depicted in Fig 1, starts with executing a suspicious PE file in a sandbox environment, ensuring isolation from external systems. This controlled execution allows for the collection of essential data, including memory features, system calls, and function calls. Subsequently, these collected data are preprocessed and used to train various ML-based algorithms, which can enhance the malware detection model.
Unlike static analysis, dynamic malware analysis requires the execution of code in a time-restricted, closed environment, which can be resource-intensive. Research endeavors, such as studies by Firdausi et al. [23] and Luckett et al. [24], have utilized system calls as key features for training traditional ML models like k-Nearest Neighbor (k-NN), Decision Tree, Support Vector Machine (SVM), and Naive Bayes. Furthermore, studies by Pirscoveanu et al. [25], and Tobiyama et al. [26] have focused on the extraction of features from API calls, evaluating the effectiveness of ML algorithms, including Random Forest, k-NN, and Convolutional Neural Networks (CNN) in dynamic malware analysis. These approaches highlight the dynamic method’s capacity for dealing with complex malware detection challenges, although it requires significant time and resource allocation.
@@ II-A3 Hybrid Analysis
A methodology integrating static and dynamic techniques, hybrid analysis, is another malware detection technique [27]. This concept has been explored in various studies. For example, Santoso et al. [28] utilized a combination of Artificial Neural Networks (ANN) and CNN for malware detection. Focusing on Android malware, Zhu et al.[29] proposed an innovative framework using the Merged Sparse Autoencoder (MSAE), which is an unsupervised learning algorithm demonstrating its effectiveness.
Adding to this, Tong and Yan [30] developed a method that combines static and dynamic analysis for mobile malware detection. This method compares system call patterns of benign and malicious applications with the dynamic analysis applied to unknown applications. Subsequent offline comparison of these pattern sets further validates the unknown application’s nature. Their results show the advantages of the hybrid approach over methods relying solely on static or dynamic analysis. Altaher and Barukab [31] also proposed a hybrid methodology for Android malware detection that leverages API calls and application permissions, further substantiating the potential of hybrid techniques in this field.
@@ II-B Online Malware Detection
Online malware detection stands out as a distinct approach in the cybersecurity domain. Unlike static, dynamic, or hybrid methods from file classification that analyze specific malware samples, online detection monitors the entire system in real-time, which enables the capture of malware at any moment, regardless of its activity level. This technique focuses on the behavior of the entire machine rather than individual malware behaviors.
Key contributions in this area include the work of Watson et al. [32], who developed a system using performance metrics to build SVM, achieving a 90% accuracy rate. Azmandian et al. [33] proposed intrusion-based detection techniques, while Abdelsalam et al. [34] introduced a sequential k-means clustering algorithm for anomaly detection, specifically designed for a standard 3-tier architecture on an OpenStack Testbed. Their approach leverages virtual machine systems and resource utilization features but shows limitations in detecting low-resource-utilization malware.
Further research was done by McDole et al. [35], who examined various CNN models to determine their suitability for malware detection in cloud Infrastructure as a Service (IaaS). Their subsequent study [36] compared the process-level performance metrics of different deep learning models in the context of online malware detection in cloud IaaS environments.
Similarly, Kimmel et al. [37, 38] presented a comprehensive analysis of the effectiveness of several ML models for online malware detection, focusing on system features describing processes in a virtual machine. They emphasized the use of CNNs, which are known for their simplicity and effective representation in 2D format. Abdelsalam et al. [39] extended this concept by employing a 3-dimensional CNN to enhance classifier accuracy and specifically target low-profile malware, achieving an accuracy rate of 90%.
@@ III Explainability in Machine Learning
Figure 2: Explainable Machine Learning Concept
DL-based models enable machines to develop complex hierarchical data patterns, which play a key role in tasks like classification or detection. These black-box models, by layering and integrating various levels of data representation, can enhance the predictive power of systems. However, this increased complexity often obscures the internal decision-making process, which can lead to questions about their decision logic [40, 41].
In contrast, white-box models offer a more transparent approach. They are designed to be easily interpretable, which allows users to understand how input data is transformed into predictions or decisions. This transparency is particularly valuable in fields where understanding the reasoning behind a decision is as important as the decision itself [42, 43].
@@
This is where the concept of XAI comes into the picture. As depicted in Fig 2, today’s AI systems typically involve training the data, undergoing the machine learning process, and providing prediction for end-users. In contrast, XAI goes a step further. It can deliver high-accuracy predictions and provides clear, justifiable explanations for these outcomes. With higher interpretability, the reasons behind AI predictions become more comprehensible to humans, which boosts the trustworthiness and reliability of the model’s predictions.
In the comprehensive study, Blanco-Justicia and Domingo-Ferrer. [44] discussed the seven characteristics that define XAI for enhancing transparency and efficacy in AI systems as follows.
@@
Hence, in the context of XAI, it is important to understand the general classification of ML-based models, which is illustrated in Fig 3. This figure presents a comprehensive taxonomy of ML models and XAI techniques and provides a clear framework for understanding this field.
@@
As depicted in Fig 3, ML models can be broadly classified into Transparent and Opaque models. Transparent models are inherently explainable. These models are straightforward enough that they do not require additional post-hoc explainability techniques, i.e., techniques provide explanations only after the training process has finished. However, as indicated by the dashed arrow in Fig 3, when these models become more complex, post-hoc explainability may still be needed to improve clarity and ensure human interpretability.
On the other hand, Opaque models, often referred to as black-box models, are characterized by their high accuracy yet present challenges in interpretation. Due to their complexity, they require the use of post-hoc explainability methods. The goal of post-hoc explainability is to make the outcomes of ML-based models more transparent, understandable, and trustworthy to humans.
Post-hoc explainability can be further divided into two types: model-agnostic and model-specific methods. Model-agnostic methods have a variety of explainability techniques and are versatile enough to be applied to any black-box model. In contrast, model-specific methods are applicable only to certain types of models and limit their utility to specific cases.